(8%) Problem 3: Vector A has a given B of A+B tri
Answer #1

a)Magnitude of B = magnitude of A * tan\theta = 6.5*tan42=5.85

b)A=6.5i

B=5.85j

A-B=6.5i-5.85j

Magnitude of A-B=sqrt(6.52+5.852)=8.74

c)

A+B=6.5i+5.85j

Magnitude of A-B=sqrt(6.52+5.852)=8.74